OpenStudy (anonymous):

What is the ratio 15 to 35 written as a fraction in lowest terms?

OpenStudy (anonymous):

@Lady.Liv1776

OpenStudy (luffingsails):

15/35 = 3/7 = lowest term.

OpenStudy (anonymous):

THANK YOU

OpenStudy (luffingsails):

Welcome, a ratio is nothing more than a fraction that you can simplify if possible.